OpenStudy (anonymous):

whats the answer to 7(d+2) - 2(2d+1) = 21

OpenStudy (anonymous):

7(d+2) - 2(2d+1) = 21 7d+14-4d-2=21 3d+12=21 3d=9 d=3

OpenStudy (anonymous):

d=3

OpenStudy (anonymous):

7 (2+d)-2 (1+2 d) = 21 7 2+7 d-2 1-2 2 d = 21 2 7+7 d-1 2-2 2 d = 21 2 7+7 d-1 2-2^(1+1) d = 21 2 7+7 d-1 2-4 d = 21 14+7 d-1 2-4 d = 21 14+7 d-2-4 d = 21 = 3 d+12 = 21